OpenStudy (accidentalaichan):

Can someone help review my answer? 3x + y = 11 –3x + 2y = 1 Okay, so 3x and -3x cancel out add y and 2y to equal 3y? 12 / 3y 4 = y?

jigglypuff314 (jigglypuff314):

Correct! :) then to get x plug what you got for y back into one of the original equations and solve :)

OpenStudy (accidentalaichan):

so, 4 + 3x = 12 12 - 4 = 8 8 / 3x?

OpenStudy (anonymous):

Ok, so if you put em together the 3x's cancel out each other so you get \[3y=12\]

jigglypuff314 (jigglypuff314):

mmm original equation was 3x + y = 11 not 3x + y = 12 right?

OpenStudy (accidentalaichan):

It's a system of equations ^^;;

OpenStudy (accidentalaichan):

you add 1 and 11 to get 12?

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

mmm but when solving for x you just plug what you got for y into one of the original equations so plug in y=4 into 3x + y = 11 and solve for x

OpenStudy (accidentalaichan):

so 3x + 4 = 11? Okay, that would be 3x = 9 then x = 3?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

:D Good Job Explaining JigglyPuff

OpenStudy (accidentalaichan):

Oh, thanks! I get it now!